OpenStudy (anonymous):

A rectangular pool has dimensions of 40 ft. and 60 ft. The pool has a patio area around it that is the same width on all sides.If the patio area equals the area of the pool, how wide is the distance from the pool edge to the patio edge? If x represents the width of the patio, then which of the following equations could be used to solve the word problem? x² + 100x + 2,400 = 0 x² - 50x + 600 = 0 x² + 50x - 600 = 0

OpenStudy (anonymous):

@gerryliyana

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

based on the text area of pool = area of patio

OpenStudy (anonymous):

and we've known that the are of the pool = 40 ft x 60 ft = 2400 ft^2

OpenStudy (anonymous):

because the area of pool = area of patio = 2400, we are allowed to write The area of the pool plus the patio = 2(2400) = 4800

OpenStudy (anonymous):

Okay..?

OpenStudy (anonymous):

The area of the pool plus the patio can also be written in (2x+40)*(2x+60), so (2x+40)*(2x+60)=4800 4x^2 + 120x + 80x + 2400 = 4800 4x^2 + 200x + 2400 = 4800 dividing both sides of the equation by 4 yields x^2 + 50x + 600 = 1200 x^2 + 50x + 600 - 1200 = 0 x^2 + 50x - 600 = 0

OpenStudy (anonymous):

@xo_kansasprincess_xo : got it?

OpenStudy (anonymous):

x² + 50x - 600 = 0 is the answer

OpenStudy (anonymous):

if you are satisfied with my answer, don't forget to fan me and take the time to write a testimonial about me, i will appreciate it!

OpenStudy (anonymous):

http://gamutech.com We'll help with questions you have there and there's a chat no lag like OS and you won't get suspended